# FUSE : A Ridge and Random Forest-Based Metric for Evaluating MT in Indigenous Languages

### v2

@@ 1 Introduction
MT has made significant advancements in recent years, largely driven by neural machine translation (NMT) models  Lyu et al. (2024). However, evaluating the quality of translations remains a major challenge, particularly for low-resource Indigenous languages.Traditional MT evaluation metrics such as Bilingual Evaluation Understudy (BLEU)  Papineni et al. (2002a) Translation Edit Rate (TER)  Snover et al. , and Character n-gram F-score (ChrF)  Popović rely on surface-level token overlap, which fails to capture semantic correctness, fluency, and linguistic structure—critical factors in evaluating translations for morphologically rich and polysynthetic languages. Indigenous languages, such as Bribri, Guarani, and Nahuatl, exhibit unique linguistic characteristics that pose challenges for conventional MT evaluation  Chen et al. (2023). These languages often lack standardized orthography, leading to multiple valid translations  Aepli et al. (2023). They feature lexical complexity, including polysynthesis and noun incorporation, which makes word segmentation and alignment with reference translations difficult  Tyers and Mishchenkova (2020). They also rely on phonetic variations, making strict token-level matching unreliable. Due to these factors, existing evaluation metrics struggle to provide reliable assessments of translation quality for Indigenous languages. While metrics such as BLEU and ChrF focus on exact token matches, they fail to account for phonetic and semantic similarities in morphologically rich languages.
@@
In this paper, we propose FUSE, a machine learning-based MT evaluation metric tailored for American Indigenous languages. The complete architecture of FUSE is illustrated in  Figure 1, showcasing its integration of lexical, phonetic, semantic, and fuzzy similarity features with hybrid regression modeling. It incorporates phonetic similarity features, addressing a critical gap in existing evaluation metrics. The model optimizes feature weighting using regression models trained on human scores, leading to improved correlation with human evaluation. We validate our metric on Spanish-to-Indigenous language translations, demonstrating superior performance over BLEU, TER, and ChrF.
@@ 3 Datasets
For our experiments, we utilize the datasets provided by the AmericasNLP 2025 Shared Task 3 on Machine Translation Metrics. This shared task focuses on the evaluation of automatic metrics for translations from Spanish into three Indigenous languages: Guarani, Bribri, and Nahuatl. Each dataset is split into training and test subsets, where the training data is used to build and tune our models, and the test set is used for final evaluation. The specific sizes of the training and test sets for each language are detailed in Table  table 1
@@ 6 Results
On the development set, Approach 5 achieves the highest overall performance, attaining the best average Spearman (0.8001) and Pearson (0.8455) correlations across all three language pairs. This variant, visualized in Figure fig. 1, employs a hybrid model combining Ridge regression (for semantic scoring) and Random Forest regression (for fluency), benefiting from the interpretability of linear models and the flexibility of ensemble-based non-linear modeling. Notably, it outperforms all other approaches on the Bribri language, likely due to its ability to capture intricate phonetic and lexical variability through feature learning. Approach 6, which replaces the fluency model with Gradient Boosting, performs comparably well—achieving top correlations for Guarani (Pearson: 0.8667) and Nahuatl (Spearman: 0.8216, Pearson: 0.8331)—suggesting that boosting methods are effective at modeling complex relationships in morphologically rich languages. In contrast, traditional feature-weighted approaches (Approaches 1–3) yield moderate results due to the absence of supervised weight optimization or the exclusive use of linear models, which limits their capacity to model non-linear dependencies. These development results are summarized in  Table 2.
On the held-out test set, a similar trend is observed: Approach 5 (RaaVa 2) achieves the highest average correlation with human annotations, ranking first in the shared task. As shown in  Table 3, this consistency across both development and test sets highlights the generalization capability of the ensemble architecture and validates the inclusion of phonetic, semantic, lexical, and fuzzy features. These findings further underscore the importance of integrating diverse linguistic signals with adaptive feature learning for MT evaluation in orthographically variable and low-resource Indigenous languages.
@@
  • Note: Best-performing scores in each column are highlighted in red and bold. Results are based on dev set correlations with human annotations.

### v3

@@ 1 Introduction
MT has made significant advancements in recent years, largely driven by neural machine translation (NMT) models  Lyu et al. (2024); Vats et al. (2025). However, evaluating the quality of translations remains a major challenge, particularly for low-resource Indigenous languages.Traditional MT evaluation metrics such as Bilingual Evaluation Understudy (BLEU)  Papineni et al. (2002a) Translation Edit Rate (TER)  Snover et al. , and Character n-gram F-score (ChrF)  Popović rely on surface-level token overlap, which fails to capture semantic correctness, fluency, and linguistic structure—critical factors in evaluating translations for morphologically rich and polysynthetic languages. Indigenous languages, such as Bribri, Guarani, and Nahuatl, exhibit unique linguistic characteristics that pose challenges for conventional MT evaluation  Chen et al. (2023); Raja and Vats (2025). These languages often lack standardized orthography, leading to multiple valid translations  Aepli et al. (2023). They feature lexical complexity, including polysynthesis and noun incorporation, which makes word segmentation and alignment with reference translations difficult  Tyers and Mishchenkova (2020). They also rely on phonetic variations, making strict token-level matching unreliable. Due to these factors, existing evaluation metrics struggle to provide reliable assessments of translation quality for Indigenous languages. While metrics such as BLEU and ChrF focus on exact token matches, they fail to account for phonetic and semantic similarities in morphologically rich languages.
@@
In this paper, we propose FUSE, a machine learning-based MT evaluation metric tailored for American Indigenous languages. The complete architecture of FUSE is illustrated in  Figure˜1, showcasing its integration of lexical, phonetic, semantic, and fuzzy similarity features with hybrid regression modeling. It incorporates phonetic similarity features, addressing a critical gap in existing evaluation metrics. The model optimizes feature weighting using regression models trained on human scores, leading to improved correlation with human evaluation. We validate our metric on Spanish-to-Indigenous language translations, demonstrating superior performance over BLEU, TER, and ChrF.
@@ 3 Datasets
For our experiments, we utilize the datasets provided by the AmericasNLP 2025 Shared Task 3 on Machine Translation Metrics. This shared task focuses on the evaluation of automatic metrics for translations from Spanish into three Indigenous languages: Guarani, Bribri, and Nahuatl. Each dataset is split into training and test subsets, where the training data is used to build and tune our models, and the test set is used for final evaluation. The specific sizes of the training and test sets for each language are detailed in Table  table˜1
@@ 6 Results
On the development set, Approach 5 achieves the highest overall performance, attaining the best average Spearman (0.8001) and Pearson (0.8455) correlations across all three language pairs. This variant, visualized in Figure fig.˜1, employs a hybrid model combining Ridge regression (for semantic scoring) and Random Forest regression (for fluency), benefiting from the interpretability of linear models and the flexibility of ensemble-based non-linear modeling. Notably, it outperforms all other approaches on the Bribri language, likely due to its ability to capture intricate phonetic and lexical variability through feature learning. Approach 6, which replaces the fluency model with Gradient Boosting, performs comparably well—achieving top correlations for Guarani (Pearson: 0.8667) and Nahuatl (Spearman: 0.8216, Pearson: 0.8331)—suggesting that boosting methods are effective at modeling complex relationships in morphologically rich languages. In contrast, traditional feature-weighted approaches (Approaches 1–3) yield moderate results due to the absence of supervised weight optimization or the exclusive use of linear models, which limits their capacity to model non-linear dependencies. These development results are summarized in  Table˜2.
On the held-out test set, a similar trend is observed: Approach 5 (RaaVa 2) achieves the highest average correlation with human annotations, ranking first in the shared task. As shown in  Table˜3, this consistency across both development and test sets highlights the generalization capability of the ensemble architecture and validates the inclusion of phonetic, semantic, lexical, and fuzzy features. These findings further underscore the importance of integrating diverse linguistic signals with adaptive feature learning for MT evaluation in orthographically variable and low-resource Indigenous languages.
@@
  • Note: Best-performing scores in each column are highlighted in green and bold. Results are based on dev set correlations with human annotations.